OpenStudy (afrodiddle):

The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function?

OpenStudy (afrodiddle):

-4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

OpenStudy (afrodiddle):

I know B is ruled out.

OpenStudy (welshfella):

work out what f(2) is

OpenStudy (welshfella):

= -3 * f(1) - 2* f(0) = ?

OpenStudy (welshfella):

= -36 + 8 = ?

OpenStudy (afrodiddle):

28

OpenStudy (welshfella):

no -28

OpenStudy (afrodiddle):

-28

OpenStudy (afrodiddle):

I know, my key did not click for some reason :P

OpenStudy (welshfella):

so you should be able to pick out the correct choice now

OpenStudy (afrodiddle):

A

OpenStudy (welshfella):

the first 3 terms are -4 , 12 and -28

OpenStudy (welshfella):

yes

OpenStudy (afrodiddle):

I'm amazing at history but I have a tough time with Algebra -.-

OpenStudy (welshfella):

yea we all have strengths and weaknesses

OpenStudy (afrodiddle):

Well :) thank you. I gave you a medal.

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

its funny with algebra - many people sudden;y 'see' it and it goes from tough to quite easy really quickly

OpenStudy (welshfella):

practice as much as you can